# Cancellation of phonon hopping in trapped ions by modulation of the trap potential

### v1

@@ II.2 Non-perturbative modulation of a harmonic potential
Lau and James [14] has shown that a phase shift which is faster and applicable to any number of phonons can be constructed by modulating the trap potential instead of irradiating the ions with laser light. We describe here this trap-potential modulation approach for a single harmonic oscillator whose Hamiltonian is given by $h_{0}=\hbar\omega_{0}(a^{\dagger}a+1/2)$. As shown in Appendix C, when we apply a quadratic potential,
@@ III Results
We demonstrate the C3PO by numerically solving the time-dependent Schrödinger equation (TDSE) from $t=0$ to $T$, where $T$ is set to the period of the 50:50 beam-splitter between the neighbouring modes. In order to solve the TDSE efficiently and accurately, the semi-global method [19] has been adopted. There are two sources of error in the C3PO; (i) the algorithmic error due to the imperfect cancellation of the phonon hopping as explained in II.1.2 and (ii) the gate error due to the finite pulse duration of $P^{\mathrm{LJ}}$. Because the algorithmic error is absent for $M=2$, we first examine the 2-mode C3PO to determine the optimal parameters which minimize the gate error. Then, both the algorithmic and the gate error are investigated for $M=3$. Hereafter, when the instantaneous $\pi$-phase shift is applied instead of $P^{\mathrm{LJ}}$, we call the DD the ideal DD. We evaluate the error by $\braket{E}\equiv|\bra{\psi_{0}}I-U\ket{\psi_{0}}|=1-|\bra{\psi_{0}}U\ket{\psi_{0}}|$ | for an initial state |$\ket{\psi_{0}}$, where $U$ represents the pulse sequence of the ideal DD or the C3PO.
@@
Figure 1: The populations of a 2-mode phonon state during the C3PO with $d=27.6\,\mathrm{\mu m}$. (a) $T_{P}=4.0\,\mathrm{\mu s}$ and (b) $T_{P}=1.0\,\mathrm{\mu s}$. The initially occupied phonon state is shown by the solid curve, while the other states are shown by the dashed curves. The application of $P^{\mathrm{LJ}}$ is indicated by the vertical shaded areas.
Figure 2: The populations of a 2-mode phonon state during the C3PO. $T_{P}=4.0\,\mathrm{\mu s}$ and $d=43.8\,\mathrm{\mu m}$. The initially occupied phonon state is shown by a solid curve while the other states are shown by dashed curves. The application of $P^{\mathrm{LJ}}$ is indicated by the vertical shaded areas.
We can reduce the error by shortening the pulse duration or/and slowing down the phonon hopping. For example, as shown in Fig. 1 (b), when $T_{P}$ is set to $1\,\mathrm{\mu s}$, the error is lowered to be $\braket{E}=2.2\times 10^{-6}$. Instead of making $T_{P}$ four times smaller, which requires an improvement in the experimental setup, we may make the phonon hopping slower by four times, which can be easily achieved by replacing $d$ by $4^{(1/3)}d$. As shown in Fig. 2, for $d=43.8\,\mathrm{\mu m}$, which corresponds to $\kappa_{1,0}/(2\pi)=0.47\,\mathrm{kHz}$, the error is lowered further to $\braket{E}=2.2\times 10^{-8}$.
@@ III.2.1 Ideal DD
Figure 3: The populations of a 3-mode phonon state during the single cycle of the ideal DD. The pulse sequecne is defined by Eq. (10). The initially occupied phonon state is shown by a solid curve while the other states are shown by dashed curves. The application of the instantaneous $\pi$-phase shifts is indicated by the vertical dotted lines.
@@
Although the pulse sequence is designed so that the initial state is recovered at $t=T$, the initial state is almost recovered at $t=T/2$ in Fig. 4 (a), and the dynamics from $t=0$ to $T/2$ is repeated from $t=T/2$ to $T$. This is because $P_{1}$ applied at $t=T/4$ flips the sign of the nearest neighbour coupling terms, $a_{1}^{\dagger}a_{0}+\mathrm{h.c.}$ and $a_{2}^{\dagger}a_{1}+\mathrm{h.c.}$, which have the dominant contribution to the dynamics, while $P_{2}$ applied at $t=T/2$ flips the sign of the non-nearest neighbour coupling term $a_{2}^{\dagger}a_{0}+\mathrm{h.c.}$, whose contribution is eight times smaller in magnitude because $\kappa_{2,0}\propto(2d)^{-3}$. This observation supports the validity of the truncation described in II.1.2, where the pulse sequence is constructed to cancel the hopping between ions in the distance smaller than or equal to $\eta d$.
The ideal repeated-DD $\tilde{U}_{\mathrm{DD}}^{n_{r}}$ derived from Eq. (21) can be defined similarly as in Eq. (II.1.3) and an example with $n_{r}=5$ is illustrated in Fig. 5 (a). We can see the five cycles of phonon hopping and its cancellation. We note that the error scaling given in II.1.3 is not tight and the actual scaling for fixed $\kappa_{1,0}$ and $N_{\mathrm{BP}}$ is estimated as $\mathcal{O}(T^{4}n_{r}^{-2})$ from the numerical simulation of the 3-mode DD. Here, the error for the repeated-DD with $n_{r}=5$ is calculated to be $\braket{E}=1.9\times 10^{-6}$, which is $n_{r}^{2}=25$ times smaller than the error obtained with $n_{r}=1$.
Figure 4: The populations of a 3-mode phonon state during the single cycle of (a) the ideal DD and (b) the C3PO. The initially occupied phonon state is shown by a solid curve while the other states are shown by dashed curves. The application of the instantaneous $\pi$-phase shifts is indicated by the vertical dotted lines in (a) while the application of $P^{\mathrm{LJ}}$ is indicated by the vertical shaded areas in (b).
Figure 5: The populations of a 3-mode phonon state during the five cycles of (a) the ideal DD and (b) the C3PO. The initially occupied phonon state is shown by a solid curve while the other states are shown by dashed curves. The application of the instantaneous $\pi$-phase shifts is indicated by the vertical dotted lines in (a) while the application of $P^{\mathrm{LJ}}$ is indicated by the vertical shaded areas in (b).
@@ III.2.2 C3PO
By replacing $P_{j}$ in Eq. (21) by $P^{\mathrm{LJ}}_{j}$, a numerical simulation of the 3-mode C3PO was performed with $n_{r}=1$ and $n_{r}=5$. The results of the simulation with $n_{r}=1$ and $n_{r}=5$ are shown in Fig. 4 (b) and Fig. 5 (b), respectively. The errors obtained for $n_{r}=1$ and $n_{r}=5$ are $\braket{E}=4.4\times 10^{-5}$ and $\braket{E}=2.6\times 10^{-6}$, respectively.
@@ IV Conclusion
We have proposed a method called C3PO to control phonon hopping among the local modes in a chain of trapped ions by applying a sequence of phase-shift operations realized by the modulation of the trap potential. The present method can be applied to a system having any number of phonons and modes, which has been demonstrated by the numerical simulations of the cancellation of the phonon hopping among local modes. By applying the C3PO to cancel the phonon hopping among all the modes but the chosen set of modes $\mathcal{S}$, we can implement the multi-mode entangling gate for $\mathcal{S}$. It should be noted that the C3PO can be implemented to the boson sampling [7] and the logical two-qubit gate for the GKP state [20]. For the next step for the experimental demonstration of the C3PO, we need to examine possible sources of error such as imperfect control of the phase shift pulse, motional heating, and the vibrational decoherence. We note that the logical two-qubit gate for the GKP states encoded using two radial modes of a single ion was recently demonstrated [21], but this approach can be applied to a system whose mode number is less than or equal to 3 because it relies on the coupling between the electronic and vibrational degrees of freedom of a single ion.
@@ Appendix A Hamiltonian
We consider $M$ trapped ions oscillating along $\tilde{x}$ axis. The Hamiltonian of $j$-th ion is given by $H_{j}=\tilde{p}_{j}^{2}/(2m)+m\tilde{\omega}_{0}^{2}\tilde{x}_{j}^{2}/2$, where $m$ is the ion’s mass and $\tilde{\omega}_{0}$ is the oscillation angular frequency. The Coulomb interaction between two cations located at positions $\bm{r}_{1}$ and $\bm{r}_{2}$ can be approximated as [20]
@@
where $\tilde{\kappa}_{j,k}=e^{2}/(4\pi\epsilon_{0}d_{j,k}^{3}m\tilde{\omega}_{0})$ and $x_{0}=\sqrt{\hbar/(m\omega_{0})}$. As can be seen in Eq. (A), the first term of the Coulomb interaction shifts the oscillation frequency while the second term mixes the different modes. Because the frequency shift depends on $j$ as represented by $\sum_{k}\tilde{\kappa}_{j,k}$, $P^{\mathrm{LJ}}$ should also depends on $j$, which induces additional complexity.
Therefore, we replace $\tilde{\omega}_{0}$ by the $j$-dependent frequency $\tilde{\omega}_{j}$ in a manner that the shifted oscillation frequency becomes $j$-independent. First, the Hamiltonian can be written as
@@
$x_{0,j}=\sqrt{\hbar/(m\tilde{\omega}_{j})}$, and $\tilde{\kappa}_{j,k}$ is redefined as $\tilde{\kappa}_{j,k}=e^{2}/(4\pi\epsilon_{0}d_{j,k}^{3}m\tilde{\omega}_{j,k})$ with $\tilde{\omega}_{j,k}=\sqrt{\tilde{\omega}_{j}\tilde{\omega}_{k}}$. We eliminate the $j$-dependence of the oscillation frequency by requiring
@@
Then, by introducing the scaled coordinate and momentum defined respectively as $x=\tilde{x}/x_{0}$ and $p=\tilde{p}/p_{0}$ with $x_{0}=\sqrt{\hbar/(m\omega_{0})}$ and $p_{0}=\sqrt{\hbar m\omega_{0}}$, the Hamiltonian is given by
@@
By using the relations, $x_{j}=(a^{\dagger}_{j}+a_{j})/\sqrt{2}$ and $p_{j}=\mathrm{i}(a^{\dagger}_{j}-a_{j})/\sqrt{2}$, the Hamiltonian in the Fock basis is obtained as
@@
Appendix B Phase shift and beam-splitter
@@
When the two-mode beam-splitter is defined as $B^{\mathrm{TM}}_{1,0}(\theta)=\exp(-\mathrm{i}\theta(a_{1}^{\dagger}a_{0}+a_{1}a_{0}^{\dagger}))$ and the phase shift is defined as $P_{1}(\phi)=\exp(-\mathrm{i}\phi a_{1}^{\dagger}a_{1})$, we derive
@@ Appendix C Time-dependent harmonic oscillator
where $\mathcal{H}_{n}$ is the Hermite polynomial and $b$ is related with $\omega(t)$ by
@@
When the boundary conditions, $b(t\leq 0)=b(t\geq T_{P})=1$ and $\omega(0)=\omega(T_{P})=\omega_{0}$, are satisfied, $\psi(t\geq T_{P})$ coincides with $\varphi_{n}$ except the $n$-dependent phase factor. In the interaction picture with respect to $h_{0}$, $\psi$ acquires the phase $\exp(-\mathrm{i}(n+1/2)\phi)$ at $t\geq T_{P}$, where $\phi$ is defined by Eq. (19).
@@
During the application of $P^{\mathrm{LJ}}$, the phonon number is not necessarily to be preserved. For example, during the application of $P^{\mathrm{LJ}}$ with $T_{P}=1.0\,\mathrm{\mu s}$ in Fig. 1 (b), the populations of the phonon-number-preserving states decrease while the states which do not preserve the phonon number are populated as shown in Fig. 7.
Figure 6: The time-dependent frequency $\omega(t)$ and the $b$ function for the long pulse ($T_{P}=4.0\,\mathrm{\mu s}$) with $T_{\mathrm{ud}}=4.4T_{0}=2.0\,\mathrm{\mu s}$ and $k=0.0529$.

### v2

@@ II.2 Non-perturbative modulation of a harmonic potential
Lau and James [14] have shown that a phase shift which is faster and applicable to any number of phonons can be constructed by modulating the trap potential instead of irradiating the ions with laser light. Here we describe this trap-potential modulation approach for a single harmonic oscillator whose Hamiltonian is given by $h_{0}=\hbar\omega_{0}(a^{\dagger}a+1/2)$. As shown in Appendix C, when we apply a quadratic potential,
@@ III Results
We demonstrate the C3PO by numerically solving the time-dependent Schrödinger equation (TDSE) from $t=0$ to $T$, where $T$ is set to the period of the 50:50 beam splitter between the neighboring modes. In order to solve the TDSE efficiently and accurately, the semiglobal method [19] has been adopted. There are two sources of error in the C3PO: (i) the algorithmic error due to the imperfect cancellation of the phonon hopping as explained in Sec. II.1.2 and (ii) the gate error due to the finite pulse duration of $P^{\mathrm{LJ}}$. Because the algorithmic error is absent for $M=2$, we first examine the two-mode C3PO to determine the optimal parameters which minimize the gate error. Then, both the algorithmic and the gate error are investigated for $M=3$. Hereafter, when the instantaneous $\pi$-phase shift is applied instead of $P^{\mathrm{LJ}}$, we call the DD the ideal DD. We evaluate the error by $\braket{E}\equiv|\bra{\psi_{0}}I-U\ket{\psi_{0}}|=1-|\bra{\psi_{0}}U\ket{\psi_{0}}|$ | for an initial state |$\ket{\psi_{0}}$, where $U$ represents the pulse sequence of the ideal DD or the C3PO.
@@
Figure 1: The populations of a two-mode phonon state during the C3PO with $d=27.6\,\mathrm{\mu m}$. (a) $T_{P}=4.0\,\mathrm{\mu s}$ and (b) $T_{P}=1.0\,\mathrm{\mu s}$. The initially occupied phonon state is shown by the solid curve, while the other states are shown by the dashed curves. The application of $P^{\mathrm{LJ}}$ is indicated by the vertical shaded areas.
Figure 2: The populations of a two-mode phonon state during the C3PO. $T_{P}=4.0\,\mathrm{\mu s}$ and $d=43.8\,\mathrm{\mu m}$. The initially occupied phonon state is shown by a solid curve, while the other states are shown by dashed curves. The application of $P^{\mathrm{LJ}}$ is indicated by the vertical shaded areas.
We can reduce the error by shortening the pulse duration or/and slowing down the phonon hopping. For example, as shown in Fig. 1(b), when $T_{P}$ is set to $1\,\mathrm{\mu s}$, the error is lowered to be $\braket{E}=2.2\times 10^{-6}$. Instead of making $T_{P}$ four times smaller, which requires an improvement in the experimental setup, we may make the phonon hopping slower by four times, which can be easily achieved by replacing $d$ with $4^{(1/3)}d$. As shown in Fig. 2, for $d=43.8\,\mathrm{\mu m}$, which corresponds to $\kappa_{1,0}/(2\pi)=0.47\,\mathrm{kHz}$, the error is lowered further to $\braket{E}=2.2\times 10^{-8}$.
@@ III.2.1 Ideal DD
Figure 3: The populations of a three-mode phonon state during the single cycle of the ideal DD. The pulse sequence is defined by Eq. (10). The initially occupied phonon state is shown by a solid curve, while the other states are shown by dashed curves. The application of the instantaneous $\pi$-phase shifts is indicated by the vertical dotted lines.
@@
Although the pulse sequence is designed so that the initial state is recovered at $t=T$, the initial state is almost recovered at $t=T/2$ in Fig. 4(a), and the dynamics from $t=0$ to $T/2$ is repeated from $t=T/2$ to $T$. This is because $P_{1}$ applied at $t=T/4$ flips the sign of the nearest-neighbor coupling terms, $a_{1}^{\dagger}a_{0}+\mathrm{h.c.}$ and $a_{2}^{\dagger}a_{1}+\mathrm{h.c.}$, which have the dominant contribution to the dynamics, while $P_{2}$ applied at $t=T/2$ flips the sign of the non-nearest-neighbor coupling term $a_{2}^{\dagger}a_{0}+\mathrm{h.c.}$, whose contribution is eight times smaller in magnitude because $\kappa_{2,0}\propto(2d)^{-3}$. This observation supports the validity of the truncation described in Sec. II.1.2, where the pulse sequence is constructed to cancel the hopping between ions in the distance smaller than or equal to $\eta d$.
The ideal repeated DD $\tilde{U}_{\mathrm{DD}}^{n_{r}}$ derived from Eq. (21) can be similarly defined as in Eq. (II.1.3) and an example with $n_{r}=5$ is illustrated in Fig. 5(a). We can see the five cycles of phonon hopping and its cancellation. We note that the error scaling given in Sec. II.1.3 is not tight and the actual scaling for fixed $\kappa_{1,0}$ and $N_{\mathrm{BP}}$ is estimated as $O(T^{4}n_{r}^{-2})$ from the numerical simulation of the three-mode DD. Here, the error for the repeated DD with $n_{r}=5$ is calculated to be $\braket{E}=1.9\times 10^{-6}$, which is about $n_{r}^{2}=25$ times smaller than the error obtained with $n_{r}=1$.
Figure 4: The populations of a three-mode phonon state during the single cycle of (a) the ideal DD and (b) the C3PO. The pulse sequence is defined by Eq. (21). The initially occupied phonon state is shown by a solid curve, while the other states are shown by dashed curves. The application of the instantaneous $\pi$-phase shifts is indicated by the vertical dotted lines in (a), while the application of $P^{\mathrm{LJ}}$ is indicated by the vertical shaded areas in (b).
Figure 5: The populations of a three-mode phonon state during the five cycles of (a) the ideal DD and (b) the C3PO. The pulse sequence is obtained by repeating $U_{\mathrm{DD}}$ defined by Eq. (21). The initially occupied phonon state is shown by a solid curve, while the other states are shown by dashed curves. The application of the instantaneous $\pi$-phase shifts is indicated by the vertical dotted lines in (a), while the application of $P^{\mathrm{LJ}}$ is indicated by the vertical shaded areas in (b).
@@ III.2.2 C3PO
By replacing $P_{j}$ in Eq. (21) by $P^{\mathrm{LJ}}_{j}$, a numerical simulation of the three-mode C3PO was performed with $n_{r}=1$ and $n_{r}=5$. The results of the simulation with $n_{r}=1$ and $n_{r}=5$ are shown in Figs. 4(b) and 5(b), respectively. The errors obtained for $n_{r}=1$ and $n_{r}=5$ are $\braket{E}=4.4\times 10^{-5}$ and $\braket{E}=2.6\times 10^{-6}$, respectively.
@@
III.2.3 Application of C3PO to the two-mode beam splitter
Here we implement the two-mode beam splitter by applying the C3PO to cancel unwanted phonon hoppings by the method described in Sec. II.1.2. We choose |$\ket{\psi_{0}}=\ket{1,1,1}$ as the initial state and consider the beam splitter between $q_{0}$ and $q_{1}$. First, we denote the subset to which the entangling operation defined by Eq. (5) is applied as $\mathcal{S}$ and the rest of the modes as $\mathcal{S}^{c}$. Then, the pulse sequence of the DD is constructed so that $P$ is not applied to $q^{\prime}\in\mathcal{S}$, where $\mathcal{S}=\{q_{0},\,q_{1}\}$ and $\mathcal{S}^{c}=\{q_{2}\}$. This is equivalent to considering the two-mode DD between $q^{\prime}$ and $q_{2}$. Therefore, we apply $P$ to $q_{2}$ at $t=T/2$ and $t=T$.
The results of the simulation with $n_{r}=1$ and $n_{r}=5$ are shown in Figs. 6 and 7, respectively. Because the input to the 50:50 beam splitter is |$\ket{1,1}$, the output should be (|$(\ket{2,0}+\ket{0,2})/\sqrt{2}$ due to the Hong–Ou–Mandel effect [20, 21]. Therefore, the final state is |$\ket{\psi_{f}}=(\ket{1,2,0}+\ket{1,0,2})/\sqrt{2}$, where both the populations of |$\ket{1,2,0}$ and |$\ket{1,0,2}$ increase and become close to 0.5 as expected. The error is defined as $\braket{E_{B}}\equiv|\bra{\psi_{f}}U_{\mathrm{ideal}}-U\ket{\psi_{0}}|=1-|\bra{\psi_{f}}U\ket{\psi_{0}}|$ |, where $U_{\mathrm{ideal}}=I_{2}\otimes B_{1,0}(T_{\mathrm{50:50}})$ is the desired unitary operator implementing the 50:50 beam splitter between $q_{1}$ and $q_{0}$, |$\ket{\psi_{f}}=U_{\mathrm{ideal}}\ket{\psi_{0}}$ is the desired final state, and $U$ represents the pulse sequence of the ideal DD or the C3PO. When $n_{r}=1$, the error is obtained to be $\braket{E_{B}}=4.6\times 10^{-2}$ for the ideal DD and $\braket{E_{B}}=4.5\times 10^{-2}$ for the C3PO, and when $n_{r}=5$, the error can be suppressed to be $\braket{E_{B}}=1.8\times 10^{-3}$ for the ideal DD and $\braket{E_{B}}=1.6\times 10^{-3}$ for the C3PO.
Figure 6: The populations of a three-mode phonon state during the two-mode beam splitter implemented by (a) the ideal DD and (b) the C3PO. The initial state |$\ket{1,1,1}$ (black) and the desired final states, |$\ket{1,2,0}$ (red) and |$\ket{1,0,2}$ (blue), are shown by solid curves, while the other states are shown by dashed curves. The application of the instantaneous $\pi$-phase shifts is indicated by the vertical dotted lines in (a), while the application of $P^{\mathrm{LJ}}$ is indicated by the vertical shaded areas in (b).
Figure 7: The populations of a three-mode phonon state during the two-mode beam splitter implemented by the five cycles of (a) the ideal DD and (b) the C3PO. The initial state |$\ket{1,1,1}$ (black) and the desired final states, |$\ket{1,2,0}$ (red) and |$\ket{1,0,2}$ (blue), are shown by solid curves, while the other states are shown by dashed curves. The application of the instantaneous $\pi$-phase shifts is indicated by the vertical dotted lines in (a), while the application of $P^{\mathrm{LJ}}$ is indicated by the vertical shaded areas in (b).
@@ IV Conclusion
We have proposed a method called C3PO to control phonon hopping among the local modes in a chain of trapped ions by applying a sequence of phase-shift operations realized by the modulation of the trap potential. The present method can be applied to a system having any number of phonons and modes, which has been demonstrated by the numerical simulations of the cancellation of the phonon hopping among local modes. By applying the C3PO to cancel the phonon hopping among all the modes except the chosen set of modes $\mathcal{S}$, we can implement the multi-mode entangling gate for $\mathcal{S}$. We have shown that the two-mode beam splitter can be implemented between a set of two modes chosen from a chain of three modes by canceling the unwanted phonon hopping by the C3PO. It should be noted that the C3PO can be applied to the boson sampling [7] and the logical two-qubit gate for the GKP state [22]. For the experimental demonstration of the C3PO, we need to examine possible sources of error such as imperfect control of the phase-shift pulse, motional heating, and vibrational decoherence. We note that the logical two-qubit gate for the GKP states encoded using two radial modes of a single ion was recently demonstrated [23], but this approach can be applied to a system composed of three or less modes because it relies on the coupling between the electronic and vibrational degrees of freedom of a single ion.
@@ Appendix A Hamiltonian
We consider $M$ trapped ions oscillating along the $\tilde{x}$ axis. The Hamiltonian of the $j$-th ion is given by $H_{j}=\tilde{p}_{j}^{2}/(2m)+m\tilde{\omega}_{0}^{2}\tilde{x}_{j}^{2}/2$, where $m$ is the ion’s mass and $\tilde{\omega}_{0}$ is the oscillation angular frequency. The Coulomb interaction between two cations located at positions $\bm{r}_{1}$ and $\bm{r}_{2}$ can be approximated as [22]
@@
where $\tilde{\kappa}_{j,k}=e^{2}/(4\pi\epsilon_{0}d_{j,k}^{3}m\tilde{\omega}_{0})$ and $x_{0}=\sqrt{\hbar/(m\omega_{0})}$. As can be seen in Eq. (A), the first term of the Coulomb interaction shifts the oscillation frequency, while the second term mixes the different modes. Because the frequency shift depends on $j$ as represented by $\sum_{k}\tilde{\kappa}_{j,k}$, $P^{\mathrm{LJ}}$ should also depends on $j$, which induces additional complexity.
Therefore, we replace $\tilde{\omega}_{0}$ by the $j$-dependent frequency $\tilde{\omega}_{j}$ in a manner that the shifted oscillation frequency becomes $j$ independent. First, the Hamiltonian can be written as
@@
$x_{0,j}=\sqrt{\hbar/(m\tilde{\omega}_{j})}$, and $\tilde{\kappa}_{j,k}$ is redefined as $\tilde{\kappa}_{j,k}=e^{2}/(4\pi\epsilon_{0}d_{j,k}^{3}m\tilde{\omega}_{j,k})$ with $\tilde{\omega}_{j,k}=\sqrt{\tilde{\omega}_{j}\tilde{\omega}_{k}}$. We eliminate the $j$ dependence of the oscillation frequency by requiring
@@
Then, by introducing the scaled coordinate and momentum defined, respectively, as $x=\tilde{x}/x_{0}$ and $p=\tilde{p}/p_{0}$ with $x_{0}=\sqrt{\hbar/(m\omega_{0})}$ and $p_{0}=\sqrt{\hbar m\omega_{0}}$, the Hamiltonian is given by
@@
By using the relations $x_{j}=(a^{\dagger}_{j}+a_{j})/\sqrt{2}$ and $p_{j}=\mathrm{i}(a^{\dagger}_{j}-a_{j})/\sqrt{2}$, the Hamiltonian in the Fock basis is obtained as
@@
Appendix B Phase shift and beam splitter
@@
When the two-mode beam splitter is defined as $B^{\mathrm{TM}}_{1,0}(\theta)=\exp(-\mathrm{i}\theta(a_{1}^{\dagger}a_{0}+a_{1}a_{0}^{\dagger}))$ and the phase shift is defined as $P_{1}(\phi)=\exp(-\mathrm{i}\phi a_{1}^{\dagger}a_{1})$, we derive
@@ Appendix C Time-dependent harmonic oscillator
where $\mathcal{H}_{n}$ is the Hermite polynomial and $b$ is related to $\omega(t)$ by
@@
When the boundary conditions, $b(t\leq 0)=b(t\geq T_{P})=1$ and $\omega(0)=\omega(T_{P})=\omega_{0}$, are satisfied, $\psi(t\geq T_{P})$ coincides with $\varphi_{n}$ except for the $n$-dependent phase factor. In the interaction picture with respect to $h_{0}$, $\psi$ acquires the phase $\exp(-\mathrm{i}(n+1/2)\phi)$ at $t\geq T_{P}$, where $\phi$ is defined by Eq. (19).
@@
During the application of $P^{\mathrm{LJ}}$, the phonon number is not necessarily preserved. For example, during the application of $P^{\mathrm{LJ}}$ with $T_{P}=1.0\,\mathrm{\mu s}$ in Fig. 1(b), the populations of the phonon-number-preserving states decrease, while the states which do not preserve the phonon number are populated as shown in Fig. 9.
Figure 8: The time-dependent frequency $\omega(t)$ and the $b$ function for the long pulse ($T_{P}=4.0\,\mathrm{\mu s}$), with $T_{\mathrm{ud}}=4.4T_{0}=2.0\,\mathrm{\mu s}$ and $k=0.0529$.